# The Cost of Dynamic Reasoning: Demystifying AI Agents and Test-Time Scaling from an AI Infrastructure Perspective

### v1

@@ I Introduction
Recent progress in large language models (LLMs) has shifted from scaling model size or pretraining data to improving inference-time behavior, a direction known as test-time scaling[48]. Test-time scaling is designed to enhance model performance by allocating additional computation during inference without modifying the model’s parameters. This includes techniques such as multi-sample decoding (e.g., Best-of-N[51], Self-Consistency [54]), step-by-step prompting (e.g., Chain-of-Thought [55], Least-to-Most Prompting [70], Self-Ask [38]), and structured search strategies (e.g., Tree-of-Thought [63], Graph-of-Thought [5]). These approaches promote more deliberate and interpretable reasoning within the LLM, enabling it not only to recognize patterns but also to derive conclusions, generate explanations, and solve tasks that require step-by-step logic.
@@
Figure 1: Overview of test‑time scaling. (a) Conventional LLMs map inputs directly to outputs in a single forward pass, with no explicit intermediate reasoning. (b) Reasoning‑enhanced LLMs internally create intermediate steps—sampling alternative responses or extending token sequences—to deepen or diversify their thought process. (c) AI agents augment this reasoning by 1 planning and invoking external tools, 2 observing the outcomes and adapting their internal reasoning accordingly, and iteratively refining their decision-making until they 3 generate the final answer.
@@
Deploying these reasoning-enhanced LLMs, however, comes at an immense computational cost. Even in current static reasoning models which follow fixed input-output mappings without external tool interaction (Figure 1(a,b)), LLMs run on thousands of GPUs, whose power, cooling, and capital costs drive monthly expenses into the tens of millions of dollars [44]. A single ChatGPT query is estimated to consume about ten times the electricity of a typical web search [10] and requires a substantial amount of cooling water [50]. As a result, hyperscalers are investing at an unprecedented scale. Meta has committed over \$10 billion to AI infrastructure [20], and Microsoft, in collaboration with OpenAI, operates custom AI supercomputers that draw tens to hundreds of megawatts. As an example, xAI’s Colossus AI supercomputer alone employs approximately 100,000 Nvidia H100 GPUs, consuming 150 megawatts in total, with the GPUs alone accounting for around 70 megawatts [59]. For perspective, traditional hyperscale data centers typically draw between 10 and 100 megawatts [15], while advanced semiconductor fabs, such as those operated by Samsung and TSMC, consume several hundred megawatts each [16, 4]. Analysts forecast that total AI infrastructure spending will surpass \$1 trillion within this decade [11], raising serious concerns about power grid sustainability, environmental impact, and the economic viability of large-scale LLM deployment.
@@
Given this landscape, the emergence of AI agents powered by LLMs with dynamic reasoning threatens to exacerbate these already formidable infrastructure pressures dramatically. Unlike static reasoning models, dynamic reasoning represents an advanced form of test-time scaling that significantly boosts capabilities through active interaction with external environments (Figure 1(c)). Specifically, AI agents continuously plan, invoke external tools, observe outcomes, and iteratively refine their reasoning, often performing dozens of inference calls to satisfy a single user request [64, 47, 68]. Without substantial system-level innovations, per-request computational costs could increase by orders of magnitude, making large-scale deployment of agents economically and environmentally prohibitive. Industry leaders are already responding to these challenges: OpenAI’s upcoming Stargate cluster is projected to consume multiple “gigawatts” of power, with costs potentially reaching \$500 billion [32]. Yet, despite these developments, the computer architecture and systems research communities have largely focused on static LLMs, leaving the infrastructure implications of dynamic reasoning workloads underexplored.
@@
To address this critical gap, this paper presents a rigorous, quantitative evaluation of the computational and infrastructural costs of dynamic reasoning. We systematically characterize resource utilization, latency implications, and energy demands inherent in the iterative execution patterns of AI agents. Our analysis highlights the critical system-level challenges faced when deploying AI agents and identifies opportunities for optimization through architectural improvements, enhanced inference algorithms, and intelligent resource allocation strategies. To the best of our knowledge, this work is the first to provide a system-level characterization of dynamic reasoning in AI agents, grounded in empirical analysis of end-to-end infrastructure behavior across diverse agentic workflows. A central contribution–and key objective–of our study is to quantitatively assess the AI infrastructure costs associated with dynamic reasoning deployments, and to inform and caution the broad research community about the urgent need for sustainable, efficient design principles to bridge the gap between advanced algorithmic capabilities and practical, scalable, and sustainable deployment.
@@ II-A Definition of AI Agents
AI agents are inference-time frameworks that extend the capabilities of large language models by enabling multi-step reasoning, adaptive decision-making, and interaction with the external environment. Unlike conventional LLM applications that produce a single output from a static prompt, AI agents operate through iterative internal reasoning and external actions at inference time. At each iteration, the agent may generate an intermediate reasoning result, call an external tool, such as a search engine, calculator, or code interpreter, and incorporate the output into its subsequent decisions. This process allows the agent to refine its strategy, retrieve missing information, and revise its reasoning dynamically in response to evolving task demands. While this adaptivity enhances the agent’s ability to handle complex and open-ended problems, it also leads to variability in the number of model calls, tool usage patterns, and overall computational cost.
@@ II-B Core Components and Workflows of AI Agents
This core reasoning capability of the AI agent is further supported by memory, plan, and tools. 2 Memory plays a critical role in enabling the agent to maintain continuity across reasoning steps. It stores short-term interaction traces, such as prior LLM outputs and tool responses, as well as long-term knowledge, including user preferences or experience from past interactions. 3 Plan organizes the agent’s objective into a sequence of subtasks or a directed acyclic graph (DAG) of interdependent actions. By maintaining an explicit plan, the agent can prioritize actions, track progress, and make forward-looking decisions that align with the overall task structure. 4 Tools extend the agent’s capabilities beyond text generation by enabling interaction with external environments. At each step, the agent analyzes its current context, generates a structured command specifying the desired tool and input, executes the tool call, and incorporates the resulting output into its context. This output is then used to guide the next stage of reasoning.
@@ II-C Test-Time Scaling in AI Agents
Test-time scaling refers to methods that improve the reasoning performance of pretrained LLMs at inference time by increasing the amount of computation used for inference without modifying model parameters [55, 63, 54, 70, 48]. Representative techniques include Chain-of-Thought [55], which guides the model to produce intermediate reasoning steps through carefully crafted prompts, and Tree-of-Thought [63], which expands the reasoning space by exploring multiple reasoning paths. Both approaches guide the model to perform step-by-step reasoning, effectively leveraging its internal reasoning capabilities without modifying the model parameters.
AI agents build upon this paradigm by implementing test-time reasoning not through prompt design alone, but through multi-step decision-making that integrates tool use and maintenance of intermediate reasoning state. Unlike conventional prompt-based methods that operate within a static input-output mapping, agents dynamically coordinate multiple model invocations and tool interactions, adapting their behavior based on intermediate outcomes. This form of dynamic reasoning enables agents to respond to new information, revise prior decisions, and handle real-time tasks involving external environments. As such, AI agents redefine test-time scaling by moving beyond conventional inference approaches that rely solely on the model’s internal reasoning abilities. Agentic workflows actively gather information, plan actions, and iteratively refine their reasoning over multiple iterations. This paradigm shift introduces new challenges in efficiency, latency, and resource management, highlighting the need for a system-level analysis of the behavior of AI agents.
@@ II-D Motivation
While AI agents have demonstrated promising capabilities through iterative reasoning and tool-augmented workflows, their growing complexity introduces profound challenges at the systems and infrastructure level. Unlike conventional single-turn LLM inference where computation is bounded to a single forward pass, agentic execution involves dynamically evolving control flows, multiple rounds of LLM inference, and external tool interactions. These behaviors incur significant compute overhead, amplify memory pressure, and introduce unpredictable latency and utilization patterns.
Despite these operational complexities, prior research on AI agents has largely focused on improving task success rates and qualitative reasoning behavior [64, 47, 68], with little attention paid to the underlying system and its deployment costs. Questions central to the deployment and scaling of such agents—such as GPU under-utilization due to agent’s adaptive control flow, memory bottlenecks from ever-growing contexts, tail-latency behavior under load, or energy implications of multi-step inference—remain largely unexamined. Consequently, existing architecture and systems optimizations for LLMs, which target static, single-pass workloads, may fall short in capturing or addressing the dynamic and iterative characteristics unique to AI agents.
This paper is motivated by the urgent need to fill this gap. To the best of our knowledge, this work is the first to present a rigorous, system-level characterization of AI agents, grounded in empirical measurement across diverse agent designs and tasks. We argue that without a principled understanding of the system-level implications of dynamic reasoning, the community risks building infrastructure optimized for yesterday’s workloads. A systems-oriented perspective is therefore critical to guide the design of sustainable, efficient, and scalable serving infrastructures. Our study takes this first step by analyzing the computational and infrastructural costs of deploying AI agents in practice, providing actionable insights for future architecture and systems co-design.
@@ III Methodology
In this section, we describe the experimental setup used to characterize AI agents. We consider a representative set of AI agents and benchmarking workloads, covering diverse agent workflows and agentic task characteristics.
@@
CoT [55] O X X X X
ReAct [64] O O X X X
Reflexion [47] O O O X X
LATS [68] O O O O X
LLMCompiler [18] O O O X O
TABLE I: Comparison of AI agents.
@@
AI agent workflows. To analyze the system-level design space of AI agents, we investigate five representative agents: Chain-of-Thought (CoT) [55], ReAct [64], Reflexion [47], Language Agent Tree Search (LATS) [68], and LLMCompiler [18]. These agents were selected to cover a wide range of reasoning strategies, tool integrations, and planning mechanisms. Table I summarizes the presence or absence of five key capabilities across each agent.
  • Reasoning. All agents considered in this study employ a reasoning mechanism. Among them, CoT operates purely through internal reasoning without the use of any external tool (Figure 3(a)). As a baseline for comparison, we also include language-only, CoT-style static reasoning approaches within the broader definition of AI agents, despite their lack of external interactions with tools.
  • Tool use. Tool use differentiates purely language-based agents from those capable of interacting with the environment. This functionality enables agents to access real-time data or perform non-linguistic operations.
@@
  • Tree search. LATS (Figure 3(d)) leverages Monte Carlo Tree Search (MCTS) [9] to simulate and expand multiple branches of reasoning and action, allowing the agent to evaluate different candidate paths before making a decision. By simulating multiple possible future paths, the agent can make more informed decisions and select optimal action sequences.
  • Structured planning. LLMCompiler incorporates a structured multi-step planning and streaming for asynchronous task execution to minimize latency. During the planning phase, LLMCompiler analyzes task dependencies and constructs a directed acyclic graph (DAG) that organizes future tool calls into an execution plan. This enables multiple dependent actions to be generated within a single LLM invocation. As the plan is constructed, intermediate tool calls are streamed to the execution stage, allowing the scheduler to overlap planning and tool calls via asynchronous execution. Together, these features can help reduce repeated reasoning and lower end-to-end latency (Figure 3(e)).
In general, we utilized the official open-source implementations provided by the original authors of these agent workflows [65, 46, 69, 17]. Each AI agent is adapted to support our evaluation framework and benchmarks. For LATS, we further optimized its implementation to support concurrent LLM inference and parallel tool invocation because the original version [69] executes these operations sequentially, aggravating end-to-end latency.
@@
HotpotQA [61] Task Multi-hop question answering
@@
WebShop [62] Task Online shopping
@@
MATH [14] Task Math problem solving
@@
HumanEval [8] Task Programming
@@
TABLE II: Description of benchmarks.
Benchmarks. We select four popular benchmarks representative of various downstream agentic tasks, whose descriptions are summarized in Table II.
HotpotQA [61] is a question-answering benchmark that assesses the agent’s ability to accurately retrieve relevant evidence to answer multi-hop knowledge-intensive questions. We provide the Wikipedia APIs [56] as tools to solve these questions. WebShop [62] is a web-shopping benchmark where agents find the best-fit item that meets the given conditions. The agent is given web navigation tools to browse WebShop. MATH [14] is a benchmark suite of mathematics problems across various domains. Agents are equipped with access to the Wolfram Alpha API [57] for solving complex equations, as well as a Python-based calculator for simple numerical computations. HumanEval [8] evaluates the programming capability of agents. In our setup, agents are equipped with a Python execution tool that allows them to validate the generated solutions by executing self-written test code. In addition to these agentic benchmarks, we utilize a non-agentic dataset, which is the ShareGPT dataset [45], to model conventional chatbot-like LLMs, characterized by single-turn LLM inference without iterative interactions with the external environments. ShareGPT contains a collection of real conversations between users and ChatGPT [29], capturing standard interactive dialogue scenarios.
@@
LLM backend. We employed the OpenAI-compatible vLLM (version 0.6.6) server as the LLM serving infrastructure, integrated with PyTorch 2.6 and CUDA 12.8. We enabled prefix caching [19], which reduces redundant computation by reusing previously computed attention states (i.e., Key-Value cache (KV cache)) for shared input prefixes across LLM requests. As we explore in Section IV, prefix caching is particularly effective for agent workloads that involve long instructions, few-shot examples, and iterative reasoning tokens in the input prompt. Unless otherwise specified, we use Llama-3.1-8B-Instruct [25] as the default backend LLM. However, to discuss the impact of model size on cost and accuracy, we also use Llama-3.1-70B-Instruct [24] in Section V.
Hardware. Experiments were conducted on Google Cloud Platform (GCP). For the 8B model, we used the a2-highgpu-1g instance type with 12 vCPUs (6 physical cores), 85GB memory, and a single NVIDIA A100 40GB GPU. For the 70B model, we used the a2-highgpu-8g instance type with 96 vCPUs (48 physical cores), 680GB memory, and 8 NVIDIA A100 40GB GPUs.

### v2

@@ I Introduction
Recent progress in large language models (LLMs) has shifted from scaling model size or pretraining data to improving inference-time behavior, a direction known as test-time scaling[snell2024scalingllmtesttimecompute, muennighoff2025s1simpletesttimescaling]. Test-time scaling is designed to enhance model performance by allocating additional computation during inference without modifying the model’s parameters. This includes techniques such as Chain-of-Thought [cot], Tree-of-Thought [tot], and others [touvron2023llama2openfoundation, self_cot, least_cot, press2023measuring, graph_of_thought]. These approaches promote more deliberate and interpretable reasoning within the LLM, enabling it not only to recognize patterns but also to derive conclusions, generate explanations, and solve tasks that require step-by-step logic.
@@
Figure 1: Overview of reasoning strategies in LLM-based systems. (a) Conventional LLMs map inputs directly to outputs in a single forward pass, with no explicit intermediate reasoning. (b) Reasoning‑enhanced LLMs internally create intermediate steps (e.g., sampling alternative responses or extending token sequences) to deepen or diversify their thought process. (c) AI agents augment this reasoning by 1 planning and invoking external tools, 2 observing the outcomes and adapting their internal reasoning, and iteratively refining their decision-making until they 3 generate the final answer.
@@
Deploying these reasoning-enhanced LLMs, however, comes at an immense computational cost. Even in current static reasoning models which follow fixed input-output mappings without external tool interaction (Figure 1(a,b)), LLMs run on thousands of GPUs, whose power, cooling, and capital costs drive monthly expenses into the tens of millions of dollars [openai_cost_2023]. A single ChatGPT query is estimated to consume about ten times the electricity of a typical web search [AI_energy_footprint] and requires a substantial amount of cooling water [AI_water_footprint]. As a result, hyperscalers are investing at an unprecedented scale. Meta has already committed over \$10 billion to AI infrastructure [meta_10_billion], and Microsoft operates custom AI supercomputers that draw tens to hundreds of megawatts. As an example, xAI’s Colossus AI supercomputer alone employs approximately 100,000 Nvidia H100 GPUs, consuming 150 megawatts in total, with the GPUs alone accounting for around 70 megawatts [xAI_collosus]. For perspective, traditional hyperscale data centers typically draw between 10 and 100 megawatts [hyperscale_datacenter], while advanced semiconductor fabs, such as those operated by Samsung and TSMC, consume several hundred megawatts each [tsmc_fab_power, fab_power]. Analysts forecast that total AI infrastructure spending will surpass \$1 trillion within this decade [capex_forecast], raising serious concerns about power grid sustainability and the economic viability of large-scale LLM deployment.
@@
Given this landscape, the emergence of AI agents powered by LLMs with dynamic reasoning threatens to exacerbate these already formidable infrastructure pressures dramatically. Unlike static reasoning models, dynamic reasoning represents an advanced form of test-time scaling that significantly boosts capabilities through active interaction with external environments (Figure 1(c)). Specifically, AI agents continuously plan, invoke external tools, observe outcomes, and iteratively refine their reasoning, often performing dozens of inference calls to satisfy a single user request [react, reflexion, lats]. Without substantial system-level innovations, per-request computational costs could increase by orders of magnitude, making large-scale deployment of agents economically and environmentally prohibitive. Industry leaders are already responding to these challenges: OpenAI’s planned Stargate project [openai_stargate] and Meta’s next generation AI data center Hyperion [meta_hyperion_prometheus] are each projected to require multiple “gigawatts” of power capacity, with costs reaching hundreds of billions of dollars. Yet, despite these developments, the computer architecture community has largely focused on static LLMs, leaving the infrastructure implications of dynamic reasoning workloads underexplored.
@@
To address this critical gap, this paper presents a rigorous, quantitative evaluation of the computational and infrastructural costs of dynamic reasoning. We systematically characterize resource utilization, latency implications, and energy demands inherent in the iterative execution patterns of AI agents. While the serving characteristics of static reasoning LLMs are well understood within the research community, our work presents the first comprehensive, system-level characterization of agent serving costs across diverse configurations and workloads. Each component of our analysis—including the characterization of agent workflows themselves, serving performance, and the impact of test-time scaling—quantifies a distinct dimension of this cost structure, collectively building a unified understanding of its infrastructure-level implications. Our analysis highlights the critical system-level challenges faced when deploying AI agents and identifies opportunities for optimization through architectural improvements, enhanced inference algorithms, and intelligent resource allocation strategies. To the best of our knowledge, this work is the first to provide a system-level characterization of dynamic reasoning in AI agents, grounded in quantitative analysis of end-to-end infrastructure behavior across diverse agentic workflows [^1] Open-sourced at https://github.com/VIA-Research/AgentBench.. A key contribution and objective of our study is to quantitatively assess the AI infrastructure cost of dynamic reasoning deployments, and to inform and caution the research community about the urgent need for sustainable, efficient design principles to bridge the gap between advanced algorithmic capabilities and practical, scalable, and sustainable deployment.
@@ II-A Definition of AI Agents
AI agents are inference-time frameworks that extend the capabilities of LLMs by enabling multi-step reasoning, adaptive decision-making, and interaction with the external environment. Unlike conventional LLM applications that produce a single output from a static prompt, AI agents operate through iterative internal reasoning and external actions at inference time. At each iteration, the agent may generate an intermediate reasoning result, call an external tool (e.g., search engine, calculator, or code interpreter), and incorporate the output into its subsequent decisions. This process allows the agent to retrieve missing information and refine its strategy dynamically in response to evolving task demands. While this adaptivity enhances the ability to handle complex and open-ended problems, it also leads to variability in the LLM calls, tool usage patterns, and overall computational cost.
@@ II-B Core Components and Workflows of AI Agents
This core reasoning capability of the AI agent is further supported by memory, plan, and tools. 2 Memory plays a critical role in enabling the agent to maintain continuity across reasoning steps by storing short-term interaction traces as well as long-term knowledge, including user preferences or experience from past interactions. 3 Plan organizes the agent’s objective into a sequence of subtasks or a directed acyclic graph (DAG) of interdependent actions. By maintaining an explicit plan, the agent can prioritize actions, track progress, and make forward-looking decisions that align with the overall task structure. 4 Tools extend the agent’s capabilities beyond text generation by enabling interaction with external environments. At each step, the agent analyzes its current context, generates a structured command specifying the desired tool and input, executes the tool call, and incorporates the resulting output into its context. This output is then used to guide the next stage of reasoning.
@@ II-C Test-Time Scaling in AI Agents
Test-time scaling refers to methods that improve the reasoning performance of pretrained LLMs at inference time by increasing the amount of computation used for inference without modifying model parameters [cot, tot, self_cot, least_cot, snell2024scalingllmtesttimecompute]. Representative techniques include Chain-of-Thought [cot], which guides the model to produce intermediate reasoning steps through carefully crafted prompts, and Tree-of-Thought [tot], which expands the reasoning space by exploring multiple reasoning paths. These approaches guide the model to perform step-by-step reasoning, effectively leveraging its internal reasoning capabilities while keeping its parameters fixed.
AI agents build upon this paradigm by implementing test-time reasoning not through prompt design alone, but through multi-step decision-making that integrates tool use and maintenance of intermediate reasoning state. Unlike conventional prompt-based methods that operate within a static input-output mapping, agents dynamically coordinate multiple model invocations and tool interactions, adapting their behavior based on intermediate outcomes. This form of dynamic reasoning enables agents to respond to new information, revise prior decisions, and handle real-time tasks involving external environments. As such, AI agents redefine test-time scaling by moving beyond conventional inference approaches that rely solely on the model’s internal reasoning abilities. This paradigm shift introduces new challenges in efficiency, latency, and resource management, highlighting the need for a system-level analysis of the behavior of AI agents.
@@ II-D Motivation
Unlike conventional single-turn LLM inference where computation is bounded to a single forward pass, agentic execution involves dynamically evolving control flows, multiple rounds of LLM inference, and external tool interactions. These behaviors introduce profound challenges at the systems and infrastructure level, incurring significant compute overhead, amplifying memory pressure, and introducing unpredictable latency and resource utilization patterns.
Despite these operational complexities, prior research on AI agents has largely focused on improving task success rates and qualitative reasoning behavior [react, reflexion, lats], with little attention paid to its deployment costs. Questions central to the deployment and scaling of such agents remain largely unexamined. Consequently, existing architecture and systems optimizations for LLMs, which target static, single-pass workloads, may fall short in capturing or addressing the dynamic and iterative characteristics unique to AI agents.
This paper is motivated by the urgent need to fill this gap. To the best of our knowledge, this work is the first to present a rigorous, system-level characterization of AI agents, grounded in quantitative measurement across diverse agent designs and tasks. We argue that without a principled understanding of the system-level implications of dynamic reasoning, the community risks building infrastructure optimized for yesterday’s workloads. A systems-oriented perspective is therefore critical to guide the design of sustainable, efficient, and scalable serving infrastructures. Our study takes this first step by analyzing the computational and infrastructural costs of deploying AI agents in practice, providing actionable insights for future architecture and systems co-design.
@@ III Methodology
Our analysis considers a representative set of AI agents and benchmarking workloads, covering diverse agent workflows and agentic task characteristics.
TABLE I: Comparison of AI agents.
@@
CoT [cot] O X X X X
ReAct [react] O O X X X
Reflexion [reflexion] O O O X X
LATS [lats] O O O O X
LLMCompiler [llmcompiler] O O O X O
@@
AI agent workflows. We investigate five representative agents: Chain-of-Thought (CoT) [cot], ReAct [react], Reflexion [reflexion], Language Agent Tree Search (LATS) [lats], and LLMCompiler [llmcompiler]. These agents were selected to cover a wide range of reasoning strategies, tool integrations, and planning mechanisms. Table I summarizes the presence or absence of five key capabilities across each agent.
  • Reasoning. All agents considered in this study employ a reasoning mechanism. Among them, CoT operates purely through internal reasoning without the use of any external tool (Figure 3(a)). As a baseline for comparison, CoT-style static reasoning approaches are considered within the broader definition of AI agents, despite their lack of external interactions with tools.
  • Tool use. Tool use differentiates purely language-based agents from those capable of interacting with the external environment. This functionality enables agents to access real-time data or perform non-linguistic operations.
@@
  • Tree search. LATS (Figure 3(d)) leverages Monte Carlo Tree Search [mcts] to simulate multiple branches of reasoning and action, allowing the agent to evaluate different candidate paths before making a decision. By simulating multiple possible future paths, the agent can make more informed decisions and select optimal action sequences.
  • Structured planning. LLMCompiler incorporates a structured multi-step planning and streaming for asynchronous task execution to minimize latency. During the planning phase, LLMCompiler analyzes task dependencies and constructs a DAG that organizes future tool calls into an execution plan. This enables multiple dependent tool calls to be generated within a single LLM invocation. As the plan is constructed, intermediate tool calls are streamed to the execution stage, allowing the scheduler to overlap planning and tool calls via asynchronous execution. Together, these features can help reduce repeated reasoning and lower end-to-end latency (Figure 3(e)).
In general, we utilized the official open-source implementations provided by the original authors of these agent workflows [react_github, reflexion_github, lats_github, llmcompiler_github]. Each AI agent is adapted to support our evaluation framework and benchmarks. For LATS, we further optimized its implementation to support concurrent LLM inference and parallel tool invocation because the original version [lats_github] executes these operations sequentially, aggravating end-to-end latency.
TABLE II: Description of benchmarks.
@@
HotpotQA [hotpotqa] Task Multi-hop question answering
@@
WebShop [webshop] Task Online shopping
@@
MATH [math] Task Math problem solving
@@
HumanEval [humaneval] Task Programming
@@
Benchmarks. We select four popular benchmarks representative of various downstream agentic tasks, whose descriptions are summarized in Table II. HotpotQA [hotpotqa] is a question-answering benchmark that assesses the agent’s ability to accurately retrieve relevant evidence to answer multi-hop knowledge-intensive questions. We provide the Wikipedia APIs [wikipedia_api] as tools to solve these questions. WebShop [webshop] is a web-shopping benchmark where agents find the best-fit item that meets the given conditions. The agent is given web navigation tools to browse WebShop. MATH [math] is a benchmark suite of mathematics problems across various domains. Agents are equipped with access to the Wolfram Alpha API [wolfram_alpha_api] for solving complex equations, as well as a Python-based calculator for simple numerical computations. HumanEval [humaneval] evaluates the programming capability of agents. In our setup, agents are equipped with a Python execution tool that allows them to validate the generated solutions by executing self-written test code. In addition to these agentic benchmarks, we utilize a non-agentic dataset, which is the ShareGPT dataset [sharegpt], to model conventional chatbot-like LLMs, characterized by single-turn LLM inference without iterative interactions with the external environments. ShareGPT contains a collection of real conversations between users and ChatGPT [chatgpt], capturing standard interactive dialogue scenarios.
@@
LLM backend. We employed the OpenAI-compatible vLLM (version 0.6.6) server as the LLM serving infrastructure, integrated with PyTorch 2.6 and CUDA 12.8. We enabled prefix caching [vllm], which reduces redundant computation by reusing previously computed attention states (i.e., Key-Value cache (KV cache)) for shared input prefixes across LLM requests. Unless explicitly stated otherwise, all experimental results are obtained with prefix caching enabled. We use Llama-3.1-8B-Instruct [llama_3.1_8b_instruct] as the default backend LLM. However, to discuss the impact of model size on cost and accuracy, we also use Llama-3.1-70B-Instruct [llama_3.1_70b_instruct] in Section V.
Hardware. Experiments were conducted on Google Cloud Platform (GCP). For the 8B model, we used the a2-highgpu-1g instance type with 12 vCPUs (6 physical cores), 85GB memory, and a single NVIDIA A100 40GB GPU [^2] In this work, we use GPU-based serving systems for our analysis because they are the de facto standard for large-scale LLM serving. Our characterization methodology and findings are architecture-agnostic and directly transferable to other accelerator platforms like Google TPUs. As detailed in the rest of this paper, key insights such as the impact of agentic control-flow serialization, long-context KV cache pressure, and idle-period underutilization are inherent to the workload characteristics of dynamic reasoning, not to any GPU/TPU-specific microarchitecture. Thus, the system-level implications we identify remain equally relevant to other AI inference accelerators, providing a foundation for future cross-architecture analyses.. For the 70B model, we used the a2-highgpu-8g instance type with 96 vCPUs (48 physical cores), 680GB memory, and 8 NVIDIA A100 40GB GPUs.